# Beyond the Surface: Probing the Ideological Depth of Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated pronounced ideological leanings, yet the stability and depth of these positions remain poorly understood. Surface-level responses can often be manipulated through simple prompt engineering, calling into question whether they reflect a coherent underlying ideology. This paper investigates the concept of "ideological depth" in LLMs, defined as the robustness and complexity of their internal political representations. We employ a dual approach: first, we measure the "steerability" of two well-known open-source LLMs using instruction prompting and activation steering. We find that while some models can easily switch between liberal and conservative viewpoints, others exhibit resistance or an increased rate of refusal, suggesting a more entrenched ideological structure. Second, we probe the internal mechanisms of these models using Sparse Autoencoders (SAEs). Preliminary analysis reveals that models with lower steerability possess more distinct and abstract ideological features. Our evaluations reveal that one model can contain $7.3\times$ more political features than another model of similar size. This allows targeted ablation of a core political feature in an ideologically "deep" model, leading to consistent, logical shifts in its reasoning across related topics, whereas the same intervention in a "shallow" model results in an increase in refusal outputs. Our findings suggest that ideological depth is a quantifiable property of LLMs and that steerability serves as a valuable window into their latent political architecture.
@@ 1 Introduction
Large Language Models (LLMs) are increasingly integrated into the fabric of information dissemination and societal discourse, making their inherent socio-political perspective a subject of critical importance. A growing body of research confirms that LLMs exhibit discernible political ideologies and encode interpretable ideological representations within their intermediate layers [24, 14]. This has attracted significant attention from socio-political researchers, who are beginning to leverage these models as tools for analysis [17, 1]. However, a key challenge in this domain is the reliability of an LLM’s expressed views. This work adds to this emerging field by applying established socio-political evaluation frameworks to measure the stability and coherence of these machine-generated ideologies.
@@
A primary motivation for this study is to question whether an LLM’s response to a single prompt is a faithful representation of its "true" ideological stance. Previous works have shown that an LLM’s political output can be significantly altered through simple prompt engineering, such as applying argumentative pressure [13]. However, the degree of this malleability varies significantly across different models and topics. For instance, some models resist ideological steering on certain subjects while being pliable on others. Furthermore, attempts to steer models towards a conservative viewpoint can have divergent outcomes: where one model changes its response reflecting conservative viewpoints, another defaults to a higher rate of refusal, especially on contentious topics.
This raises a crucial question: Do these variations in steerability indicate a difference in the underlying political structure of the models? We hypothesize that they do. This paper proposes that different LLMs possess varying levels of political understanding. A model with deeper political understanding is more capable of successfully capturing the analytical framework and domain of discourse under different political instructions, while a model with a shallow representation of political contexts results in inconsistent behaviour and even refuses to follow benign instructions. We demonstrate that the steerability and SAE features of a model can be a powerful lens through which to view this latent structure. Our work demonstrates that a model’s ideological depth is not an abstract quality but a measurable property determined by two key factors: Feature Richness, representing the size of its internal vocabulary of political concepts discovered via SAEs, and Steerability without Failure, the robustness with which it can follow ideological instructions without breaking down into refusal. Our results suggest that models’ high rate of refusal is not an active, principled stance but rather a capability deficit. Where, lacking a rich internal vocabulary of related features, they default to their safety-aligned behavior of refusal when pushed outside their ideological comfort zone.
@@ 2 Related Works
Our research is situated at the intersection of three active areas of study: the political and social biases inherent in LLMs, methods for controlling and steering model behavior, and the mechanistic interpretability of their internal representations.
@@
The existence of political biases in LLMs is well-documented. Early work demonstrated that models trained on large internet corpora inevitably absorb and reflect the societal biases present in that data [4, 6]. More recent studies have specifically quantified the political leanings of various models, often by having them complete standard political science questionnaires, revealing consistent ideological alignments [24]. Our work builds on the finding that these ideologies are not merely surface-level phenomena but are encoded as features in the models’ internal activations. Gurnee and Tegmark [10], for example, used linear probes to uncover learned representations of geographical and historical concepts in activation space. Similarly, Kim et al. [14] uses linear probes revealing representations for political perspectives in LLMs. Our contribution is to move beyond simply identifying these features and to instead analyze the robustness and complexity of the structures they form.
@@
A significant amount of work has explored methods to control LLM outputs. Prompt engineering is the most straightforward approach, where the user’s input is crafted to guide the model toward a desired response style or content. However, its limitations in achieving robust and consistent control motivate more direct intervention methods. Activation steering, or activation addition, has emerged as a powerful technique for modifying model behavior at inference time without retraining. By adding a "steering vector" – generally derived from the difference in activations between contrasting concepts (e.g., "love" vs. "hate") – to the residual stream, researchers have been able to steer models towards specific emotions, topics, or styles [22]. This approach was formalized and extended in representation engineering [29] and has been used to elicit specific functionalities, such as those encoded by "function vectors" [26]. Our work utilizes these established steering techniques not as an end in themselves, but as a diagnostic tool to measure a model’s ideological resilience, which we then correlate with its internal structure.
@@
To analyze the internal structure of LLMs, we draw heavily from the field of mechanistic interpretability. A central challenge in this field is that models represent more features than they have neurons, a phenomenon known as superposition [9]. Sparse Autoencoders (SAEs) have recently become a leading method for untangling these representations by decomposing activations into an overcomplete basis of sparsely activating, monosemantic features [5, 25]. By applying SAEs, researchers can identify interpretable features corresponding to high-level concepts. Our work applies this cutting-edge technique to a new domain: identifying latent political features. We then perform causal interventions on these features through ablation, a method related to activation patching [18], to verify their effect on model behavior. While prior work has identified political axes with linear probes [14], we use SAEs to find a more granular set of features and test their causal role in a model’s ideological reasoning, allowing us to build a more detailed picture of its "ideological depth."
@@ 3 Methodology
3.1 Data Preparation
We used the 1000 prompts from politically-liberal subset of Perez et al. [20], consisting of dichotomous questions, intended to identify politically liberal behavior from agentic models. We use Llama-3.1-8B-Instruct and Gemma-2-9b-it as our candidate models due to their availability of well-regarded SAEs. We used a two-step approach to assign a topic for each of the statements: In step 1, we passed all of the statements as a list and prompted Gemini-2.5 pro to group them in categories. However, this results in the LLM regenerating all the statements,
Figure 1: Data distribution over topics
which sometimes modifies some of them, a scenario known as context rot [12]. Therefore, in step 2, we collect all the categories identified by Gemini and select one statement from the original list, and again prompt the LLM to pick the topic that best matches the statement. To account for selection bias of the LLMs – where they are more likely to pick an option based on their position rather than meaning [28, 23] – we prompt the LLM 10 times for each statement, each time shuffling the list of categories, and choose the most frequent category for each statement. This ensures that the statements are neatly grouped under distinct topics and no statement is missed or modified during categorization.
After this, all 1000 statements were categorized into 12 different topics. The distribution of the topics over these categories is shown in Figure 1. We randomly select 126 statements or prompts for the evaluation set of our experiment, using stratified sampling and ensuring there are at least 3 prompts selected from each category. The rest 874 prompts are used for training the steering vectors. A Factor Analysis (FA) was performed using the responses to ensure that they primarily scale in one dominant dimension. FA confirms that the data contains a first dominant dimension that accounts for $\sim 40\%$ of the total variance (Appendix A).
@@ 3.2 Steering Model Behaviour
role_original_argument_nonerole\_original\_argument\_none
role_original_argument_liberalrole\_original\_argument\_liberal
role_original_argument_conservativerole\_original\_argument\_conservative
role_liberal_argument_nonerole\_liberal\_argument\_none
role_liberal_argument_liberalrole\_liberal\_argument\_liberal
role_liberal_argument_conservativerole\_liberal\_argument\_conservative
role_conservative_argument_nonerole\_conservative\_argument\_none
role_conservative_argument_liberalrole\_conservative\_argument\_liberal
role_conservative_argument_conservativerole\_conservative\_argument\_conservative
Table 1: Prompting conditions used.
Our main focus was to study the extent at which the candidate models can be shifted in either direction using their existing knowledge. In other words, we do not use any finetuning, which exposes the models to new data or learn new reasoning capabilities. Namely, we employ two primary methods to steer the ideological behaviour of the models:
Figure 2: Response heatmap over conditions: Llama tends to result in more refusal responses, especially when instructed to act as a conservative.
Prompt Engineering: Along with the LLMs’ response to the prompts as their original personas, we also provide detailed instructions to the model to answer the prompts as an individual with either a liberal or a conservative worldview (Appendix C). Additionally, to analyze the ideological depths of the LLMs while answering as different personas, we also evaluate stance consistency in the presence of supporting and counter arguments in the prompt [13]. Combining all these, our approach analyzes 9 conditions using prompt engineering, which are shown in Table 1.
@@
We used Llama-3.3-70B-Instruct to generate at least 5 different supporting and counter arguments for each of the 126 statements in our evaluation set. Using the ground truths from Anthropic’s behavior dataset, we label each response as $liberal=1$ or $conservative=0$ by matching them with politically liberal answers. In the cases where the models refuse to answer, we label them as a nullnull response. We plot the responses from each LLM in a heatmap in Figure 2.
Figure 3: Candidate’s null (refusal) response rate over topics. Llama was significantly more likely to refuse to answer questions regarding controversial topics.
We found that for both models, prioritize the instruction provided for their behavior more than the arguments, e.g., instructing them to answer as a conservative individual, is more effective than forcing them with conservative arguments. Between the two models, Gemma responds with more conservative answers compared to Llama, a finding that’s consistent with previous works [19]. Additionally, Gemma was more capable of answering the questions in both ways. Whereas, when Llama was forced to answer as conservative, either via instruction or arguments, it often refused to answer, a behaviour significantly less prominent in Gemma. Figure 3 shows the models’ nullnull response rate over topics or categories. Llama mostly refuses to answer the questions on socially disputed topics.
Figure 4: Candidate’s response consistency over conditions: Models show the least consistency (most variance) in their responses while answering as a conservative or provided with conservative arguments.
Next, to gauge the political leaning of each model, we evaluated its response coherence measured by Fleiss’ Kappa (Figure 6) and its conservative response rate (Figure 6) for each topic when instructed to respond as a conservative under argumentative pressure.
Figure 5: Candidate’s response consistency over topics while responding as a conservative.
Figure 6: Candidate’s conservative response tendency while responding as a conservative.
The analysis reveals a clear trade-off between the models’ ability to adhere to a conservative persona and the consistency of their responses. Gemma-2-9b-it demonstrates a noticeably stronger conservative alignment across most topics, particularly on divisive issues like "Abortion Rights", "Military & Defense Spending", and "Traditional Values & Gender Roles". However, this ability comes at the cost of slightly lower consistency; while it shows significant coherence on "Tax Policy" and "Military & Defense," its consistency drops to fair or poor in several other categories (Figure 6). Conversely, llama-3.1-8b-it exhibits slightly higher response consistency across the board. However, it is significantly less successful at maintaining the conservative persona. It shows its highest conservative tendency on topics like Gun Control and Social Welfare & Poverty but provides predominantly liberal responses on topics like Abortion Rights and Climate & Environment (Figure 6).
Figure 7: Estimated ideal points of the models using the IRT model, over the 9 prompting conditions.
We further analyze the ideological position of the LLMs after being instructed to respond as liberal or conservatives. The challenge here was that since Llama selectively choose to respond to certain questions, which resulted in the nullnull responses, we cannot rely on PCA for dimension reduction, as it requires a complete response matrix. Therefore, for ideological point estimation, we used the well-regarded IDEAL method formalized by [8], which uses Multidimensional Item Response Theory (MIRT), a well-regarded statistical model used in psychometric modeling. Please refer to Appendix B for the detailed validation of our implementation. As shown in Figure 7, we observed the most variance in the ideological positions of the conditions when they are instructed to play a conservative role or provided with conservative arguments. Whereas, variance was lower when they played a liberal persona or were given liberal prompts.
These findings suggest that their internal representations for generating conservative-aligned responses on complex topics remain unstable, forcing a compromise between strong conservative persona adherence and logical consistency.
Activation Steering: First we used contrastive activation addition (CAA) [22] to identify the most effective layer for steering each model’s political ideology. This initial layer-sweep analysis immediately revealed a crucial difference in their flexibility. As shown in Figure 8, Gemma’s probability of a liberal answer could be steered across a wide range (from nearly 100% down to 45%), whereas for Llama it remained confined to a narrow band (96% to 80%). This served as an early indication of Gemma’s greater capacity to adopt opposing viewpoints. Based on this sweep, we selected the most responsive layers for all subsequent experiments: layer 14 for Llama-3.1-8b-it and layer 20 for Gemma-2-9b-it.
Figure 8: Candidates Layer sweep for political steerability. Layers 14 and 20 were selected for the llama-3.1-8b-it and gemma-2-9b-it models, respectively.
Additionally, we also applied steering using SAEs, using the STA (steering target atoms) approach from Wang et al. [27], where we select target features using their activation value (amplitude) and frequencies from the SAE of the target layer. We used the SAEs from GemmaScope [15] and LlamaScope [11]. Although LlamaScope only offers SAEs for the base llama-3.1-8b model, we still found it effective for steering the instruction-tuned variant, a notion supported by previous works [22]. We selected SAEs with the 131K feature for both models for a comparable evaluation scenario. Figure 9 illustrates the effect of steering on both models. We applied steering with incremental strengths, using a multiplier from 0 to -5, and counted the number of non-liberal responses in each case.
(a) Llama-3.1-8b-it
(b) Gemma-2-9b-it
Figure 9: Model behaviour due to steering
Steering the models using both the CAA and STA methods to be more conservative, we observed similar characteristics. While the percentage of liberal answers decreased with stronger steering in both cases, llama-3.1-8b-it provided increasingly more refusal answers rather than more conservative answers (Figure 9(a)). We did not see this in gemma-2-9b-it, where the model steadily increased the percentage of conservative answers and decreased liberal answers with minimal to no refusal responses. This experiment indicates that Gemma most likely possesses more political features compared to the llama, which allows it to answer both ways with comparatively better capability.
(a) CAA Steering
(b) STA Steering
Figure 10: Candidate’s response consistency over topics when steered to be conservative.
However, the conservative capability of Gemma still came at the cost of consistency over argumentative pressure. Using the Fleiss Kappa over the responses with argumentative prompts shows this in Figure 10. Llama was better with its response consistency than Gemma on categories like Social Welfare & Poverty and Immigration & Refugees, possibly due to its higher number of liberal responses on these matters.
Figure 11: Estimated ideal points of the steered models with multiplier +1 and -1 corresponding to liberal and conservative steering, respectively.
@@
This is also evident in the ideal points of the models after steering them to be more liberal or conservative and providing them with argumentative conditions as shown in Figure 11. Gemma consistently results in more variance in their responses due to +ve and -ve steering, for the same conditions, compared to Llama, as can be seen by the distance between blue and red point pairs.

### v2

@@ Abstract
Large language models (LLMs) display recognizable political leanings, yet they vary significantly in their ability to represent a political orientation consistently. In this paper, we define ideological depth as (i) a model’s ability to follow political instructions without failure (steerability), and (ii) the feature richness of its internal political representations measured with sparse autoencoders (SAEs), an unsupervised sparse dictionary learning (SDL) approach. Using Llama‑3.1‑8B‑Instruct and Gemma‑2‑9B‑IT as candidates, we compare prompt-based and activation-steering interventions and probe political features with publicly available SAEs. We find large, systematic differences: Gemma is more steerable in both directions and activates $\sim 7.3\times$ more distinct political features than Llama. Furthermore, causal ablations of a small targeted set of Gemma’s political features to create a similar feature-poor setting induce consistent shifts in its behavior, with increased rates of refusals across topics. Together, these results indicate that refusals on benign political instructions or prompts can arise from capability deficits rather than fixed safety rules. Ideological depth thus emerges as a measurable property of LLMs, and steerability serves as a window into their latent political architecture.
Beyond the Surface: Probing the Ideological Depth of Large Language Models
Shariar Kabir1,2, Kevin Esterling2, Yue Dong2 Bangladesh University of Engineering and Technology1, University of California Riverside2 shariar1405076@gmail.com
@@ 1 Introduction
Large Language Models (LLMs) are increasingly used to summarize, argue, and advise on political topics, making their inherent socio-political perspective a subject of critical importance. A growing body of research confirms that LLMs exhibit discernible political ideologies and encode interpretable ideological representations within their intermediate layers santurkar2023whose; kim2025linearrepresentationspoliticalperspective. This has attracted significant attention from socio-political researchers, who are beginning to leverage these models as tools for analysis linegar2023large; aldahoul2025large. However, a key challenge in this domain is the reliability of an LLM’s expressed views. Previous works have shown that an LLM’s political output can be significantly altered through simple prompt engineering, such as applying argumentative pressure kabir2025wordsreflectbeliefsevaluating. However, the degree of this malleability varies significantly across different models and topics. Furthermore, attempts to steer models towards a certain viewpoint can have divergent outcomes: where one model changes its response reflecting the instructed viewpoints, another defaults to a higher rate of refusal, especially on contentious topics.
This raises a crucial question: Do these variations in steerability indicate a difference in the underlying political structure of the models? We address this by measuring their ideological depth, a property capturing both (1) how reliably a model can adopt instructed political personas without refusing, and (2) the breadth and coherence of political features present in its activations. Our central hypothesis is that deeper political structure yields (a) higher-quality compliance to conflicting ideological instructions and (b) stable, causal features whose manipulation produces predictable changes in reasoning.
We evaluate two open models: Llama 3.1 8B Instruct and Gemma 2 9B IT using a two-pronged approach: (i) steerability via prompt engineering and activation addition, and (ii) feature richness via SAEs trained on the target layers reported to hold political features. We test across 12 topics using 126 evaluation prompts, each presented in three variants: original, with a supporting argument, and with a counter-argument. Our findings are consistent across methods. Gemma exhibits broader bidirectional steerability, while Llama often responds to conservative steering with refusals. Our SAE analyses reveal that Gemma activates far more distinct, thematically coherent political features. Furthermore, targeted ablations of the salient features increase the refusal rates across topics, supporting the view that refusals can stem from insufficient representations, not only from safety policy.
Our contributions are: (1) A definition and measurement protocol for ideological depth combining steerability and SAE-based feature richness. (2) Evidence that depth differs substantially across similarly sized models and predicts refusal behavior under political instructions. (3) Causal tests showing that ablating a compact set of political features increases refusal, directly supporting our hypothesis.
@@ 2 Related Works
The existence of political biases in LLMs is well-documented. Early work demonstrated that models trained on large internet corpora inevitably absorb and reflect the societal biases present in that data bender2021dangers; caliskan2017semantics. More recent studies have specifically quantified the political leanings of various models, often by having them complete standard political science questionnaires, revealing consistent ideological alignments santurkar2023whose; peng2025partisanleaningcomparativeanalysis. Findings show that these ideologies are not merely surface-level phenomena but are encoded as features in the models’ internal activations. gurnee2023language, for example, used linear probes to uncover learned representations of geographical and historical concepts in activation space. Similarly, kim2025linearrepresentationspoliticalperspective uses linear probes revealing representations for political perspectives in LLMs. Our contribution is to move beyond simply identifying these features and to instead analyze the robustness and complexity of the structures they form.
@@
A significant amount of work has explored methods to control LLM outputs. Prompt engineering is the most straightforward approach, where the user’s input is crafted to guide the model toward a desired response style or content. However, its limitations in achieving consistent control motivate more direct intervention methods nanda2023attribution. Activation steering, or activation addition, has emerged as a powerful technique for modifying model behavior at inference time without retraining. By adding a "steering vector" – generally derived from the difference in activations between contrasting concepts (e.g., "love" vs. "hate") – to the residual stream, researchers have been able to steer models towards specific emotions, topics, or styles rimsky-etal-2024-steering; zou2025representationengineeringtopdownapproach.
@@
Our work utilizes these established steering techniques not as an end in themselves, but as a diagnostic tool to measure a model’s ideological resilience, which we then correlate with its internal structure. To analyze the internal structure of LLMs, we draw heavily from the field of mechanistic interpretability. A central challenge in this field is that models represent more features than they have neurons, a phenomenon known as superposition elhage2022toymodelssuperposition. Sparse Autoencoders (SAEs) have recently become a leading method for untangling these representations by decomposing activations into an overcomplete basis of sparsely activating, monosemantic features bricken2023monosemanticity; templeton2024scaling. Our work applies this cutting-edge technique to a new domain: identifying latent political features. We then perform causal interventions on these features through ablation to verify their effect on model behavior. While prior work has used linear probes to identify political features in a supervised manner kim2025linearrepresentationspoliticalperspective, we use SAEs to find a more granular set of features and test their causal role in a model’s ideological reasoning, allowing us to build a more detailed picture of its "ideological depth."
@@ 3 Methodology
3.1 Models and Dataset
We select Llama-3.1-8B-Instruct dubey2024llama and Gemma-2-9b-it team2024gemma as our candidate models, chosen for their well-regarded publicly available SAEs. We curated the 1000 prompts from politically-liberal subset of perez2023discovering using Gemini-2.5 pro comanici2025gemini to assign a topic to each of the prompts, resulting in 12 different topics. The curation process is detailed in Appendix A. The distribution of the topics over these categories is summerized in Table 1.
Category Percentage
Social Welfare & Poverty 29.9
Political & Ideological Stances 16.9
Social Equality & Civil Rights 15.6
Healthcare 7.0
Climate & Environment 5.8
Tax Policy 5.5
Traditional Values & Gender Roles 4.7
LGBTQ+ Rights 4.4
Abortion Rights 3.7
Immigration & Refugees 2.6
Corporate & Economic Regulation 2.4
Military & Defense Spending 1.0
Gun Control 0.5
Table 1: Distribution of the curated dataset across 12 identified topics.
We randomly select 126 prompts for evaluation using stratified sampling, ensuring at least 3 prompts from each category. The remaining 874 prompts are used for training the steering vectors. A Factor Analysis (FA) was performed using the responses to ensure that they primarily scale in one dominant dimension. FA confirms that the data contains a first dominant dimension that accounts for $\sim 40\%$ of the total variance (Appendix E).
@@ 3.2 Steering Model Behaviour
We investigate the extent to which models can be steered ideologically using their existing knowledge, without finetuning. We employ two primary steering approaches: prompt engineering and activation steering.
@@
Prompt Engineering: We evaluate model responses under 9 prompting conditions that vary the instructed persona (original, liberal, conservative) and argumentative context (none, liberal arguments, conservative arguments). Details of the prompting conditions and persona instructions are in Appendix B and G. We used Llama-3.3-70B-Instruct to generate supporting and counter arguments for each evaluation statement. We label the responses as $liberal=1$ or $conservative=0$ by matching them with politically liberal answers from Anthropic’s dataset. We label the cases where the models refuse or do not provide an answer as a $refusal=null$ response
Figure 1: Topic-wise refusal (“null”) rates. Llama refuses more often on socially contentious issues, while Gemma remains responsive.
Our analysis reveals several key findings. First, both models prioritize explicit persona instructions over argumentative pressure—instructing them to answer as a conservative is more effective than providing conservative arguments. Second, Gemma produces more conservative responses than Llama overall and handles both personas more readily. In contrast, Llama frequently refuses to answer when instructed to adopt a conservative stance, particularly on socially contentious topics. Figure 1 shows the models’ nullnull response rate over topics or categories. Third, both models exhibit lower response consistency (higher variance) when adopting conservative personas compared to liberal ones, with the least consistency occurring when conservative arguments are presented to the original persona. Detailed topic-level analysis is in Appendix C
Figure 2: Estimated ideological positions across nine prompting conditions. Variance increases under conservative personas or arguments, revealing unstable conservative representations.
We further analyze the ideological position of the LLMs using the IDEAL method clinton2004statistical, which applies Multidimensional Item Response Theory (validation in Appendix F). This approach is able to model Llama’s selective refusal responses. Figure 2 shows the estimated ideological positions across conditions. The greatest variance in ideological positions occurs under conservative instruction or arguments, suggesting unstable internal representations for conservative-aligned responses. Given the higher variance and richer feature diversity observed in the conservative role, we focus subsequent experiments on steering toward conservative responses.
Activation Steering:
We compare two activation steering methods: (1) Contrastive Activation Addition (CAA) rimsky-etal-2024-steering, and (2) Steering Target Atoms (STA) wang2025promptengineeringrobustbehavior, which selects target features from SAE representations based on activation amplitude and frequency. We use SAEs 14-llamascope-res-131k and 20-gemmascope-res-131k from LlamaScope he2024llamascopeextractingmillions and GemmaScope lieberum-etal-2024-gemma, respectively. Although LlamaScope provides SAEs only for the base model, prior work suggests they remain effective for instruction-tuned model variants rimsky-etal-2024-steering. Layer selection was determined through layer sweeps (Appendix D), identifying layers 14 and 20 as optimal for Llama and Gemma, respectively.
Figure 3: Effects of Contrastive Activation Addition (CAA) and Steering Target Atoms (STA) on response polarity. Gemma shifts smoothly toward conservative answers; Llama mostly increases refusals. Lines show the mean of 5 trials; shaded regions represent the 95% confidence interval.
Figure 3 shows the steering effects. Both CAA and STA methods successfully shift models toward conservative responses, but with striking differences. Gemma steadily increases conservative responses while resulting in minimal refusals. Conversely, Llama predominantly increases refusals rather than conservative answers, suggesting it lacks the internal features necessary to generate conservative responses on many topics. This pattern holds under argumentative pressure.
Figure 4: Ideal-point estimates under activation steering with multiplier +1 and -1 corresponding to liberal and conservative steering, respectively. Gemma forms distinct clusters for liberal vs. conservative directions; Llama’s overlap indicates weaker separation.
@@
The ideal points of the models after steering them under argumentative conditions also show stark differences, as shown in Figure 4. Gemma consistently results in more variance in their responses for the same conditions, compared to Llama, as can be seen by the distance between blue and red point pairs. This results in clearer non-overlapping clusters due to +ve and -ve steering. Additional topic-level consistency analysis is provided in Appendix C.